The demand for a product, in dollars, is p = 2000-0.2x-0.01x2 find the consumer surplus when the sales level is 250

1
Expert's answer
2022-05-03T16:42:06-0400

Since the number of products sold is "x=250," the coresponding price is


"P=2000-0.2(250)-0.01(250)^2=1325"

The consumer surplus is as follows


"\displaystyle\int_{0}^{250}(p(x)-P)dx"

"=\displaystyle\int_{0}^{250}(2000-0.2x-0.01x^2-1325)dx"

"=[675x-0.1x^2-\dfrac{0.01x^3}{3}]\begin{matrix}
 250 \\
 0
\end{matrix}"

"=675(250)-0.1(250)^2-\dfrac{0.01(250)^3}{3}"

"=\$110416.67"